# On a generalisation of Cameron’s base size conjecture

### v2

@@ 3.2. Conjugacy classes
Now [22, 3.1.14] implies that two elements in ${\rm GL}_{n}(q)$ are conjugate if and only if they have the same Jordan decomposition on $V$, and the uniqueness of the lift $\hat{x}$ implies that we again have a complete description of unipotent classes in ${\rm PGL}_{n}(q)$. There are similar descriptions in the unitary, symplectic, and orthogonal cases and we direct the reader to [22, Sections 3.2.3, 3.3.3, 3.4.3, 3.4.4, 3.5.3, 3.5.4] for more information.
The main difference in the symplectic and orthogonal case worth highlighting, is when $p=2$, in which case all unipotent elements have Jordan decomposition $J_{2}^{s},J_{1}^{n-2s}]$ and there are at least two $G_{0}$-classes of elements with this decomposition when $s$ is even.
@@
We record [16, 3.50] that will be useful to us when dealing with field, graph and graph-field automorphisms.
@@
Suppose that $G_{0}\neq{\rm P\Omega}^{+}_{8}(q)$, let $H$ be a subgroup of $G$ and suppose that $x\in H\setminus{\rm PGL}(V)$ has prime order. Then $x^{G}\cap H\subseteq\widetilde{H}x$, where $\widetilde{H}=H\cap G_{I}$.
@@ 3.3. Probabilistic methods
Let $G$ be a finite almost simple group with socle $G_{0}$, and let $\mathscr{C}$ be the set of $G_{0}$-conjugacy classes of prime order elements in $G$. For $t\in\mathbb{R}$, we define:
@@
Results on the asymptotic behaviour of $\eta_{G}(t)$ can be extracted from [61, 1.10]. Note that the function considered in [61] is rather different to $\eta_{G}(t)$. In particular, Liebeck and Shalev sum over all $G$-conjugacy classes, whilst we are summing over all prime order $G_{0}$-classes. However, as we explain below, it is still possible to extract asymptotic results for our purposes:
@@
We have $\sum_{i=1}^{r}|x_{i}^{G}|^{-t}\leqslant\eta_{G}(t)$ for any almost simple group $G$ with socle $G_{0}$ and $t\in(0,1)$, where $x_{1},\ldots,x_{r}$ is a set of representatives of $G$-conjugacy classes of prime order elements in $G$.
@@
Note that the proof of Lemma 3.11 does not rely at all on the assumption that the elements we are summing over have prime order. In fact, by carefully inspecting [61, 1.10], we note that Liebeck and Shalev use the exact same observation in their proof and work in terms of $G_{0}$-classes. This allows us to extract the following asymptotic result:
@@
Let $G=G(q)$ be an almost simple group with socle $G_{0}$ over $\mathbb{F}_{q}$. Then the following hold:
@@
    For a fixed real number $t>0$, there exists some integer $r(t)$, such that for all almost simple groups $G=G(q)$ with rank $r\geqslant r(t)$ we have
@@
We will prove (i). The argument for (ii) is identical. If $\mathscr{S}$ denotes the set of $G_{0}$-conjugacy classes in $G$, then inspecting [61, 1.10] gives
@@
Let $G$ be a finite almost simple classical group of rank $n$ with socle $G_{0}$. If $n\geqslant 6$, then $T_{G}<1/3$. Moreover, if $G_{0}={\rm PSL}_{10}^{\epsilon}(q)$, then $T_{G}<1/5$ and if $G_{0}={\rm P\Omega}_{n}^{\epsilon}(q)$ and $n\geqslant 12$, then $T_{G}<4/15$.
@@
We now record a result on fixed point ratios of almost simple classical groups, which is an immediate consequence of [55, Theorem 1]. Note that the original result concerns all the almost simple classical groups, but we will only need the result for groups with socle ${\rm PSL}_{2}(q)$.
@@
Let $G$ be an almost simple group with socle ${\rm PSL}_{2}(q)$, where $q=p^{f}$, $H$ be a core-free subgroup of $G$, and $x\in G$ an element of prime order $r$. Then
@@
Sufficient upper bounds on $|x^{G_{I}}|$ (or $|x^{G_{0}}|$ if $x$ is a field, graph, or graph-field automorphism) are established in [16, Sections 3.3-3.5]. Combining those with [22, 3.2.8, 3.2.9, 3.2.14, 3.2.15, 3.3.10, 3.3.12, 3.3.17, 3.4.12, 3.4.15, 3.4.16, 3.5.20], which describe when prime order $G_{I}$-classes split into $G_{0}$-classes, we verify the entries in Table 5. The argument is similar in most cases, so we only treat a few cases that are slightly more involved here.
First assume that $G_{0}={\rm PSL}_{3}^{\epsilon}(q)$ and $x\in G_{I}$. Note that [22, 3.2.8, 3.3.10] imply that $|x^{G_{0}}|=|x^{G_{I}}|$, unless $x$ is unipotent with Jordan decomposition $J_{3}]$, in which case $|x^{G_{0}}|\geqslant\frac{1}{3}|x^{G_{I}}|$. Therefore, if $x$ is unipotent, then it follows from [16, 3.22] that $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{4}$. Moreover, if $s$ denotes the codimension of the largest eigenspace of the lift $\hat{x}$ of $x$ in ${\rm GL}(V)$, then we note that $s=2$ unless $x$ has Jordan form $J_{2},J_{1}]$, and so with this exception we obtain the better bound
$|x^{G_{0}}|>\frac{1}{6}\left(\frac{q}{q+1}\right)q^{6}.$
Next suppose that $x$ is semisimple. If $C_{\bar{G_{I}}}(x)$ is connected, where $\bar{G_{I}}$ denotes the ambient algebraic group of $G$, then the bounds in Table 5 are verified via [16, 3.36]. On the other hand, if $C_{\bar{G_{I}}}(x)$ is non-connected, then [15, 3.35] implies that $r$ divides $3$, and so $r=3$, and moreover
@@
Now assume that $G_{0}={\rm PSp}_{4}(q)$ and $x\in G_{I}$. Appealing to [16, 3.21] we find that $|x^{G_{0}}|>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{4}$ if $x$ is unipotent, and in particular $|x^{G_{0}}|>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{6}$ if $x$ does not have Jordan decomposition $J_{2},J_{1}^{2}]$ on $V$.
Similarly, if $x$ is semisimple and $s$ denotes the codimension of the largest eigenspace of the lift $\hat{x}$ of $x$ on ${\rm Sp}_{4}(q)$, then we note that $s>1$, and so [16, 3.36-3.37] immediately gives us
$|x^{G_{0}}|=|x^{G_{I}}|>\frac{1}{4}q^{4}$
and unless $x$ is a $t_{1}$ or $t_{1}^{\prime}$ involution, then we obtain the better bound $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$ via [16, 3.30, 3.37].
@@ 4. Fixed point ratios for classical groups
In this section we improve the fixed point ratio bounds obtained in [14, 15, 16, 17] for some low-dimensional classical groups. The goal here is to obtain sharper bounds that will allow us to use the $\eta$ function machinery discussed in Definition 3.10 and Propositions 3.13 and 3.14 in some low-dimensional cases, where a more direct approach was taken in [18] to derive results on base sizes.
Recall that $G$ is a finite almost simple classical group with socle $G_{0}$ over $\mathbb{F}_{q}$, where $q=p^{f}$ for some prime $p$. Moreover, $V$ denotes the natural module of $G_{0}$ and we will write $H$ for a non-standard subgroup of $G$. We will improve the fixed point ratio bounds in Theorem 3.9 only for certain families of classical groups. The following list contains the possible socle types of the groups we consider and we will treat each possibility in turn:
$\mathcal{B}=\{{\rm PSp}_{8}(q),{\rm PSL}^{\epsilon}_{8}(q),{\rm P\Omega}_{10}^{\epsilon}(q),{\rm PSL}_{6}^{\epsilon}(q),{\rm PSp}_{6}(q),{\rm\Omega}_{7}(q),{\rm PSL}_{5}^{\epsilon}(q),{\rm PSL}_{4}^{\epsilon}(q),{\rm PSL}_{3}^{\epsilon}(q)\}$
@@
If $G_{0}\in\mathcal{B}$, then ${\rm fpr}(x,G/H)<|x^{G_{0}}|^{-\beta}$, where $\beta$ is recorded in Table 6.
@@
We note that our bounds have a slightly weaker form than the ones given by Burness in [14, 15, 16, 17]. In particular, he bounds ${\rm fpr}(x,G/H)$ in terms of the $G$-conjugacy class of $x$, whilst we obtain an estimate in terms of the $G_{0}$-class of $x$. However, note that our different definition of $\eta_{G}$ allows us to work with this more relaxed bound.
@@
    In the majority of cases we first verify that the claim holds using Magma for small values of $q$, where our general approach is not applicable, using the methods described in Section 2.2.
@@
    ${\rm fpr}(x,G/H)\leqslant\frac{|x^{G_{I}}\cap H|}{|x^{G_{0}}|}$
    so our first step is to derive an upper bound on $|x^{G_{I}}\cap H|$ and a lower bound on $|x^{G_{0}}|$, which we either do directly, or we inspect the proofs of [14, 15, 16, 17], and we use the bounds derived there when those are sufficient.
@@
If $H$ is not of type ${\rm Sp}_{4}(q^{2})$ or ${\rm Sp}_{4}(q)\wr S_{2}$, then the claim follows by Theorem 3.9, so it remains to consider those two cases. The claim can easily be checked using Magma for $q=3$ using the function CheckFpr() provided in Section 2.2, so for the remainder of the proof we may assume that $q\geqslant 4$.
We first consider the case where $x$ is semisimple. First suppose that $H$ is of type ${\rm Sp}_{4}(q^{2})$. Let $i$ be minimal such that $r$ divides $q^{i}-1$ and $i_{0}$ be minimal such that $r$ divides $q^{2i_{0}}-1$, and note that $i_{0}=i/2$ if $i$ is even and $i_{0}=i$ otherwise. Note that we may assume that $x\in H$, since otherwise ${\rm fpr}(x,G/H)=0$. In this case $r$ must divide $|H|$, and so we have $i\in\{1,2,4,8\}$. We need to consider each possibility of $i$ in turn. The calculations for each value of $i$ are straightforward and similar, and so we will only consider the cases where $i=4$ and $i=1$ here and leave the rest for the reader to check.
First suppose that $i=4$. In this case $i_{0}=2$ and $C_{G}(x)$ is of type ${\rm GU}_{2}(q^{2})$, ${\rm GU}_{1}(q^{2})^{2}$, or ${\rm GU}_{1}(q^{2})\times{\rm Sp}_{4}(q)$. If $C_{G}(x)$ is of type ${\rm GU}_{2}(q^{2})$, then we compute
@@
If $C_{G}(x)$ is of type ${\rm GU}_{1}(q^{2})^{2}$, then we have
@@
and so ${\rm fpr}(x,G/H)<|x^{G}|^{-1/3}$ here as well.
Finally, if $C_{G}(x)$ is of type ${\rm GU}_{1}(q^{2})\times{\rm Sp}_{4}(q)$, then we compute
@@
Now suppose that $H$ is of type ${\rm Sp}_{4}(q)\wr S_{2}$ and again let $i$ be minimal such that $r|q^{i}-1$. We may again assume that $x\in H$, and in particular that $r$ divides $|H|$, since otherwise we have ${\rm fpr}(x,G/H)=0$. It follows that $i\in\{1,2,4\}$. The computations are straightforward here as well, and so, to avoid unnecessary repetition, we only consider the case $i=2$ and leave the remaining cases to the reader.
If $i=2$, then there are $11$ possibilities for $C_{G}(x)$ up to isomorphism. More precisely, $C_{G}(x)$ is one of the following types:
@@
Let $\widehat{H}={\rm Sp}_{4}(q)\wr S_{2}$ denote the lift of $H$ in ${\rm GSp}_{8}(q)$ and $V_{1}\oplus V_{2}$ denote the decomposition of $V$ into $4$-dimensional $\mathbb{F}_{q}$-spaces $V_{1},V_{2}$ preserved by $\hat{H}$. Note that since $r$ is odd, we must have $\hat{x}\in\widehat{B}={\rm Sp}_{4}(q)^{2}$. Now there are two $\widehat{H}-$classes of elements that fuse to the ${\rm GSp}_{8}(q)$-class of $\hat{x}$, namely one with representative $\hat{y},\hat{y}$], where $\hat{y}\in{\rm Sp}_{4}(q)$ is ${\rm Sp}_{4}(q)$-conjugate to $\Lambda_{1},\Lambda_{2}]$ and the one with representative $\hat{z},\hat{w}$], where $\hat{z}\in{\rm Sp}_{4}(q)$ is ${\rm Sp}_{4}(q)$-conjugate to $\Lambda_{1}^{2}]$ and $\hat{w}\in{\rm Sp}_{4}(q)$ is ${\rm Sp}_{4}(q)$-conjugate to $\Lambda_{2}^{2}]$. It now remains to compute the sizes of the $\widehat{H}-$classes of $\hat{y},\hat{y}$] and $\hat{z},\hat{w}$] to obtain a bound on $|x^{G_{I}}\cap H|$. It is easy to see that the $\widehat{H}-$class of $\hat{y},\hat{y}$] consists of all elements of the form $\hat{a},\hat{b}$], with $\hat{a},\hat{b}$ $\widehat{H}-$conjugate to $\hat{y}$, and so
$|[\hat{y},\hat{y}]^{\widehat{H}}|=\frac{|{\rm Sp}_{4}(q)|^{2}}{|{\rm GU}_{1}(q)|^{4}}.$
On the other hand, if $\hat{a},\hat{b}$] is $\widehat{H}-$conjugate to $\hat{z},\hat{w}$], then since $\widehat{H}$ can permute $V_{1}$ and $V_{2}$ either $\hat{a}\in\hat{z}^{\widehat{H}}$ and $\hat{b}\in\hat{w}^{\widehat{H}}$, or $\hat{a}\in\hat{w}^{\widehat{H}}$ and $\hat{b}\in\hat{z}^{\widehat{H}}$, and so
$|[\hat{z},\hat{w}]^{\widehat{H}}|=2\left(\frac{|{\rm Sp}_{4}(q)|}{|{\rm GU}_{2}(q)|}\right)^{2}.$
@@
Now suppose that $x$ is a unipotent element of odd prime order. If $H$ of type ${\rm Sp}_{4}(q^{2})$, then the possible Jordan forms on $V$ for an element $x\in H$ are $J_{4}^{2}],[J_{2}^{4}]$, and $J_{2}^{2},J_{1}^{4}]$. We can easily compute $|x^{G_{I}}\cap H|$ and $|x^{G_{0}}|$ using the information in [22, Table B.7] in every case and deduce that the desired bound holds. For example, if $x$ has Jordan form $J_{2}^{2},J_{1}^{4}]$, then we have
@@
and those bounds are sufficient.
@@
We now consider the case where $x$ is an involution. If $H$ is of type ${\rm Sp}_{4}(q)\wr S_{2}$, then one can easily check that the desired bound holds by inspecting the proof of [16, 2.8, 2.9], so we now assume that $H$ is of type ${\rm Sp}_{4}(q^{2})$.
If $x$ is not semisimple with $C_{G}(x)$ of type ${\rm Sp}_{4}(q)^{2}$, then the bounds found in the proof of [16, 3.2] are sufficient, so we may now assume that $q$ is odd and $C_{G}(x)$ is of type ${\rm Sp}_{4}(q)^{2}$. By [22, Section 3.4.2.2] we have $x\in G_{0}$. Now [47, 4.3.10] tells us that

### v3

@@ 3.2. Conjugacy classes
Now [22, 3.1.14] implies that two elements in ${\rm GL}_{n}(q)$ are conjugate if and only if they have the same Jordan decomposition on $V$, and the uniqueness of the lift $\hat{x}$ implies that we again have a complete description of unipotent classes in ${\rm PGL}_{n}(q)$.
There are similar descriptions in the unitary, symplectic, and orthogonal cases and we direct the reader to [22] for detailed information. It is worth noting that in the symplectic and orthogonal cases the Jordan decomposition of $x$ does not determine the $G_{I}$-class of $x$ in general. Let us briefly look at the symplectic case:
If $p$ is odd, then the $G_{0}$-classes of unipotent elements are in one-to-one correspondence with the signed partitions of the form
$(p^{a_{p}},\epsilon_{p-1}(p-1)^{a_{p-1}},(p-2)^{a_{p-2}},\ldots,\epsilon_{2}2^{a_{2}},1^{a_{1}}),$
where $\sum_{i}ia_{i}=n$, $\epsilon_{2i}=\pm$ and $a_{i}$ is even when $i$ is odd.
If $p=2$, all unipotent elements have Jordan decomposition $J_{2}^{s},J_{1}^{n-2s}]$. For each odd $s\in\{1,\ldots,n/2\}$ we get a unique $G_{I}$-class with Jordan decomposition $J_{2}^{s},J_{1}^{n-2s}]$, which we denote by $b_{s}$, and for each even $s\in\{1,\ldots,n/2\}$, we get two distinct $G_{I}$-classes with Jordan decomposition $J_{2}^{s},J_{1}^{n-2s}]$, which we label by $a_{s}$ and $c_{s}$. The orthogonal case is similar.
@@
We record [15, 3.50] that will be useful to us when dealing with field, graph and graph-field automorphisms.
@@
Suppose that $G_{0}\neq{\rm P\Omega}^{+}_{8}(q)$, let $H$ be a subgroup of $G$ and suppose that $x\in H\setminus{\rm PGL}(V)$ has prime order. Then $x^{G}\cap H\subseteq H_{I}x$, where $H_{I}=H\cap G_{I}$.
@@ 3.3. Probabilistic methods
Let $G$ be a finite almost simple classical group with socle $G_{0}$, and let $\mathscr{C}$ be the set of $G_{0}$-conjugacy classes of prime order elements in $G$. For $t\in\mathbb{R}$, we define:
@@
Results on the asymptotic behaviour of $\eta_{G}(t)$ can be extracted from [62, 1.10]. Note that the function considered in [62] is rather different to $\eta_{G}(t)$. In particular, Liebeck and Shalev sum over all $G$-conjugacy classes, whilst we are summing over all prime order $G_{0}$-classes. However, as we explain below, it is still possible to extract asymptotic results for our purposes:
@@
We have $\sum_{i=1}^{r}|x_{i}^{G}|^{-t}\leqslant\eta_{G}(t)$ for any almost simple classical group $G$ with socle $G_{0}$ and $t\in(0,1)$, where $x_{1},\ldots,x_{r}$ is a set of representatives of $G$-conjugacy classes of prime order elements in $G$.
@@
Note that the proof of Lemma 3.11 does not rely at all on the assumption that the elements we are summing over have prime order. In fact, by carefully inspecting [62, 1.10], we note that Liebeck and Shalev use the exact same observation in their proof and work in terms of $G_{0}$-classes. This allows us to extract the following asymptotic result:
@@
Let $G=G(q)$ be an almost simple classical group with socle $G_{0}$ over $\mathbb{F}_{q}$. Then the following hold:
@@
    For a fixed real number $t>0$, there exists some integer $r(t)$, such that for all almost simple classical groups $G=G(q)$ with dimension $r\geqslant r(t)$ we have
@@
We will prove (i). The argument for (ii) is identical. If $\mathscr{S}$ denotes the set of $G_{0}$-conjugacy classes in $G$, then inspecting [62, 1.10] gives
@@
Let $G$ be a finite almost simple classical group of dimension $n$ with socle $G_{0}$. Then the following hold:
  • (i)
    If $n\geqslant 6$, then $T_{G}<1/3$;
  • (ii)
    If $G_{0}={\rm PSL}_{10}^{\epsilon}(q)$, then $T_{G}<1/5$;
  • (iii)
    If $G_{0}={\rm P\Omega}_{n}^{\epsilon}(q)$ and $n\geqslant 12$, then $T_{G}<4/15$.
@@
We now record a result on fixed point ratios of almost simple classical groups, which is an immediate consequence of [56, Theorem 1]. Note that the original result concerns all the almost simple classical groups, but we will only need the result for groups with socle ${\rm PSL}_{2}(q)$.
@@
Let $G$ be an almost simple group with socle ${\rm PSL}_{2}(q)$ for $q\geqslant 23$, $H$ be a non-standard subgroup of $G$, and $x\in G$ an element of prime order $r$. Then
@@
Lower bounds on $|x^{G_{I}}|$ (or $|x^{G_{0}}|$ if $x$ is a field, graph, or graph-field automorphism) are established in [16, Sections 3.3-3.5]. Combining these with results in [22], which describe when prime order $G_{I}$-classes split into $G_{0}$-classes, we verify the entries in Table 6. The argument is similar in most cases, so we only treat a few cases that are slightly more involved here. In particular, we will discuss the cases where $G_{0}={\rm PSL}_{3}^{\epsilon}(q)$ and $x\in G_{I}$, or $x$ is a field or graph-field automorphism and $G_{0}={\rm PSp}_{4}(q)$ and $x\in G_{I}$.
First assume that $G_{0}={\rm PSL}_{3}^{\epsilon}(q)$ and $x\in G_{I}$. If $x$ is unipotent, then either $x=[J_{3}]$ and [16, 3.22] gives $|x^{G_{I}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$, or $x=[J_{2},J_{1}]$ and  [16, 3.22] gives $|x^{G_{I}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{4}$. Moreover, [22, 3.2.8, 3.3.10] imply that $x^{G_{I}}=x^{G_{0}}$ if $x=[J_{2},J_{1}]$, whilst $x^{G_{I}}$ splits into at most 33 $G_{0}$-classes if $x=[J_{3}]$, so the bounds in Table 6 are satisfied.
Next suppose that $x$ is semisimple and let $c$ be as defined in (1). If $c\geqslant 2$, or $c=1$ and $r\neq 3$, then the bounds in Table 6 are verified via [15, 3.36]. On the other hand, if $c=1$ and $r=3$, then [15, 3.35] gives
@@
Next assume that $G_{0}={\rm PSp}_{4}(q)$ and $x\in G_{I}$. Suppose that $x$ is unipotent. If $p$ is odd, then appealing to [15, 3.21] we find that $|x^{G_{0}}|>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{4}$, and in particular $|x^{G_{0}}|>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{6}$ if $x$ does not have Jordan decomposition $J_{2},J_{1}^{2}]$ on $V$. On the other hand, if $p=2$, then inspecting [15, Table 3.5] tells us that $|x^{G_{0}}|>\frac{1}{2}q^{4}$ if $x$ is an $a_{2}$ or $b_{1}$-involution and $|x^{G_{0}}|>\frac{1}{2}q^{6}$ if $x$ is a $c_{2}$-involution.
Similarly, if $x$ is semisimple and $s$ denotes the codimension of the largest eigenspace of the lift $\hat{x}$ of $x$ on ${\rm Sp}_{4}(q)$, then we note that $s>1$, and so [15, 3.36-3.37] immediately give
$|x^{G_{0}}|=|x^{G_{I}}|>\frac{1}{4}q^{4}.$
Moreover, unless $x$ is a $t_{1}$ or $t_{1}^{\prime}$ involution, then we obtain the better bound $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$ via [15, 3.30, 3.37].
@@ 4. Fixed point ratios for classical groups
In this section we improve the fixed point ratio bounds obtained in [14, 15, 16, 17] for some low-dimensional classical groups. The goal here is to obtain sharper bounds that will allow us to use the eta function machinery discussed in Definition 3.10 and Propositions 3.13 and 3.14 in some low-dimensional cases, where a more direct approach was taken in [18] to derive results on base sizes.
In this section, we will write $G$ for a finite almost simple classical group with socle $G_{0}$ over $\mathbb{F}_{q}$, where $q=p^{f}$ for some prime $p$. Moreover, $V$ will denote the natural module of $G_{0}$ and we will write $H$ for a non-standard subgroup of $G$ and $G_{I}$ for ${\rm Inndiag}(G_{0})$. Finally, we will denote the standard Frobenius morphism defined by $\lambda\mapsto\lambda^{q}$ by $\sigma$. We will improve the fixed point ratio bounds in Theorem 3.9 only for certain families of classical groups. The following list contains the possible socle types of the groups we consider and we will treat each possibility in turn:
$\mathcal{B}=\{{\rm P\Omega}_{10}^{\epsilon}(q),{\rm PSL}^{\epsilon}_{8}(q),{\rm PSp}_{8}(q),{\rm\Omega}_{7}(q),{\rm PSL}_{6}^{\epsilon}(q),{\rm PSp}_{6}(q),{\rm PSL}_{5}^{\epsilon}(q),{\rm PSL}_{4}^{\epsilon}(q),{\rm PSL}_{3}^{\epsilon}(q)\}$
@@
If $G_{0}\in\mathcal{B}$, and $x\in G$ has prime order $r$, then ${\rm fpr}(x,G/H)<|x^{G_{0}}|^{-\beta}$, where $\beta$ is recorded in Table 7.
@@
We note that our bounds have a slightly weaker form than the ones given by Burness in [14, 15, 16, 17]. In particular, he bounds ${\rm fpr}(x,G/H)$ in terms of the $G$-conjugacy class of $x$, whilst we obtain an estimate in terms of the $G_{0}$-class of $x$. However, note that our different definition of $\eta_{G}$ allows us to work with this slightly looser bound.
@@
    In the majority of cases we first verify that the claim holds using Magma for small values of $q$, where our general approach is not applicable. To do this, we use the methods described in Section 2.2.
@@
    ${\rm fpr}(x,G/H)\leqslant\frac{|x^{G_{I}}\cap H|}{|x^{G_{0}}|}.$
    Therefore, our first step is to derive an upper bound on $|x^{G_{I}}\cap H|$ and a lower bound on $|x^{G_{0}}|$, which we either do directly, or we inspect the proofs of [14, 15, 16, 17], and we use the bounds derived there when those are sufficient.
@@
If $H$ is not of type ${\rm Sp}_{4}(q^{2})$ or ${\rm Sp}_{4}(q)\wr S_{2}$, then the claim follows by Theorem 3.9, so it remains to consider those two cases. The claim can easily be checked using Magma for $q\leqslant 3$ using the function MaxFpr() provided in Section 2.2, so for the remainder of the proof we may assume that $q\geqslant 4$. As mentioned in Section 2.2, the computation for $q=3$ is more time-consuming than other cases we handle using Magma, but one can get output in less than an hour.
Case 1. $x$ is semisimple and $r>2$: First suppose that $H$ is of type ${\rm Sp}_{4}(q^{2})$. Let $i$ be minimal such that $r$ divides $q^{i}-1$ and $i_{0}$ be minimal such that $r$ divides $q^{2i_{0}}-1$, and note that $i_{0}=i/2$ if $i$ is even and $i_{0}=i$ otherwise. Note that we may assume that $x\in H$, since otherwise ${\rm fpr}(x,G/H)=0$. In this case $r$ must divide $|H|$, and so we have $i\in\{1,2,4,8\}$. We need to consider each possibility of $i$ in turn. The calculations for each value of $i$ are straightforward and similar, and so we will only consider the cases where $i=4$ and $i=1$ here and leave the rest for the reader to check.
First suppose that $i=4$. In this case $i_{0}=2$ and $C_{G}(x)$ is of type ${\rm GU}_{2}(q^{2})$, ${\rm GU}_{1}(q^{2})^{2}$, or ${\rm GU}_{1}(q^{2})\times{\rm Sp}_{4}(q)$. If $C_{G}(x)$ is of type ${\rm GU}_{2}(q^{2})$, then we may write $x=[\Lambda^{2}]$ for some $\sigma$-orbit $\Lambda$ and using Lemma 3.6 we compute
@@
If $C_{G}(x)$ is of type ${\rm GU}_{1}(q^{2})^{2}$, then $x=[\Lambda_{1},\Lambda_{2}]$ for some $\sigma$-orbits $\Lambda_{1},\Lambda_{2}$, and we have
@@
and so ${\rm fpr}(x,G/H)<|x^{G_{0}}|^{-1/3}$ here as well.
Finally, if $C_{G}(x)$ is of type ${\rm GU}_{1}(q^{2})\times{\rm Sp}_{4}(q)$, then $x=[\Lambda,I_{4}]$ for some $\sigma$-orbit $\Lambda$, and again using Lemma 3.6 we compute
@@
Now suppose that $H$ is of type ${\rm Sp}_{4}(q)\wr S_{2}$ and again let $i$ be minimal such that $r|q^{i}-1$. We may again assume that $x\in H$, and in particular that $r$ divides $|H|$. It follows that $i\in\{1,2,4\}$. The computations are straightforward here as well, and so, to avoid unnecessary repetition, we only consider the case $i=2$ and leave the remaining cases to the reader.
If $i=2$, then there are at most $11$ possibilities for $C_{G}(x)$ up to isomorphism. More precisely, $C_{G}(x)$ is one of the following types:
@@
Let $\widehat{H}={\rm Sp}_{4}(q)\wr S_{2}$ denote the lift of $H$ in ${\rm Sp}_{8}(q)$ and $V_{1}\oplus V_{2}$ denote the decomposition of $V$ into $4$-dimensional $\mathbb{F}_{q}$-spaces $V_{1},V_{2}$ preserved by $\hat{H}$. Note that since $r$ is odd, we must have $\hat{x}\in\widehat{B}={\rm Sp}_{4}(q)^{2}$. Now there are two $\widehat{H}-$classes of elements that fuse to the ${\rm Sp}_{8}(q)$-class of $\hat{x}$, namely one with representative $Y^{2}]$, where $Y\in{\rm Sp}_{4}(q)$ is ${\rm Sp}_{4}(q)$-conjugate to $\Lambda_{1},\Lambda_{2}]$ and the one with representative $[Z,W]$, where $Z\in{\rm Sp}_{4}(q)$ is ${\rm Sp}_{4}(q)$-conjugate to $\Lambda_{1}^{2}]$ and $W\in{\rm Sp}_{4}(q)$ is ${\rm Sp}_{4}(q)$-conjugate to $\Lambda_{2}^{2}]$. It now remains to compute the sizes of the $\widehat{H}-$classes of $Y^{2}]$ and $[Z,W]$ to obtain a bound on $|x^{G_{I}}\cap H|$. It is easy to see that the $\widehat{H}-$class of $Y^{2}]$ consists of all elements of the form $[A,B]$, with $A,B$ $\widehat{H}-$conjugate to $Y$, and so
$|[Y^{2}]^{\widehat{H}}|=\frac{|{\rm Sp}_{4}(q)|^{2}}{|{\rm GU}_{1}(q)|^{4}}.$
On the other hand, if $[A,B]$ is $\widehat{H}-$conjugate to $[Z,W]$, then since $\widehat{H}$ can permute $V_{1}$ and $V_{2}$ either $A\in Z^{\widehat{H}}$ and $B\in W^{\widehat{H}}$, or $A\in W^{\widehat{H}}$ and $B\in Z^{\widehat{H}}$, and so
$|[Z,W]^{\widehat{H}}|=2\left(\frac{|{\rm Sp}_{4}(q)|}{|{\rm GU}_{2}(q)|}\right)^{2}.$
@@
Case 2. $x$ is unipotent and $p$ is odd: If $H$ of type ${\rm Sp}_{4}(q^{2})$, then [22, 5.3.2(i)] implies that the possible Jordan forms on $V$ for an element $x\in H$ are $J_{4}^{2}],[J_{2}^{4}]$, and $J_{2}^{2},J_{1}^{4}]$. We can easily compute $|x^{G_{I}}\cap H|$ and $|x^{G_{0}}|$ using the information in [22, Table B.7] in every case and deduce that the desired bound holds. For example, suppose that $x$ has Jordan form $J_{2}^{2},J_{1}^{4}]$, and let $W$ denote the natural module of ${\rm Sp}_{4}(q^{2})$. Appealing to [22, 5.3.2(i)], we deduce that $x$ has Jordan form $J_{2},J_{1}^{2}]$ on $W$. Moreover, we note that there are two signed partitions corresponding to this Jordan decomposition, so we compute
@@
and these bounds are sufficient.
@@
Case 3. $x$ is an involution: If $H$ is of type ${\rm Sp}_{4}(q)\wr S_{2}$, then one can easily check that the desired bound holds by inspecting the proof of [16, 2.8, 2.9], so we now assume that $H$ is of type ${\rm Sp}_{4}(q^{2})$.
If $x$ is not semisimple with $C_{G}(x)$ of type ${\rm Sp}_{4}(q)^{2}$, then the bounds found in the proof of [16, 3.2] are sufficient, so we may now assume that $q$ is odd and $C_{G}(x)$ is of type ${\rm Sp}_{4}(q)^{2}$. By [22, Section 3.4.2.2] we have $x\in G_{0}$. Now [48, 4.3.10] tells us that